The test statistic of z=2.37 is obtained when testing the claim that p>0.2.

a. Identify the hypothesis test as being​ two-tailed, left-tailed, or​ right-tailed.

b. Find the​ P-value.

c. Using a significance level of α=0.10​, should we reject H0 or should we fail to reject H0​?
Answer #1

a) right-tailed

b) p-value = P(Z > 2.37) = 1 - P(Z < 2.37) = 1 - 0.9911 = 0.0089

c) since p-value is less than significance level, we should reject H0.

we reject H0